# M3: A Modular M orld Model over Streams of Tokens

### v2

@@ Abstract
Token-based world models emerged as a promising modular framework, modeling dynamics over token streams while optimizing tokenization separately. While successful in visual environments with discrete actions (e.g., Atari games), their broader applicability remains uncertain. In this paper, we introduce $\text{M}^{\text{3}}$, a modular world model that extends this framework, enabling flexible combinations of observation and action modalities through independent modality-specific components. $\text{M}^{\text{3}}$ integrates several improvements from existing literature to enhance agent performance. Through extensive empirical evaluation across diverse benchmarks, $\text{M}^{\text{3}}$ achieves state-of-the-art sample efficiency for planning-free world models. Notably, among these methods, it is the first to reach a human-level median score on Atari 100K, with superhuman performance on 13 games. Our code and model weights are publicly available at https://github.com/leor-c/M3.
Machine Learning, ICML, Deep Learning, Deep RL, World Models
@@ 1 Introduction
Figure 1: Results overview. $\text{M}^{\text{3}}$ exhibits state-of-the-art sample-efficiency performance for planning-free methods across all three benchmarks. $\dagger$ (Robine et al., 2023), $\ddagger$ (Dedieu et al., 2025).
Modeling the dynamics of the world has been a long-standing and compelling topic in reinforcement learning, with foundational ideas proposed over 30 years ago (Sutton, 1991; Schmidhuber, 1991, 2010). It has been leveraged to enhance planning performance (Schrittwieser et al., 2019), produce fictitious data for policy training (Hafner et al., 2023), or simulate interactive experiences (DeepMind, 2024; Decart et al., 2024; Agarwal et al., 2025).
@@
In this study, we focus on token-based world models (TBWMs) (Micheli et al., 2023, 2024; Cohen et al., 2024), sample-efficient reinforcement learning (RL) methods that learn the dynamics entirely within a learned token space. TBWMs offer a clear modularity between the representation and the dynamics by separating their optimization. Such modularity provides a clean and efficient way to unify different observation modalities within a shared token space, offloading representation learning to the tokenizer design step. More generally, modular systems are easier to scale, develop, evaluate, and deploy, as individual modules can be treated independently and are easier to master through divide and conquer.
While TBWMs hold promise as a general framework for handling different modalities, existing work primarily focuses on image observations and discrete actions, such as Atari games. It remains unclear whether token-based world models are truly effective for handling diverse modalities and how to design the model to achieve it. This limits the broader adoption of token-based models.
To address these limitations, we propose a modular world model, termed $\text{M}^{\text{3}}$, which implements a modular framework for general-purpose token-based world models. Specifically, $\text{M}^{\text{3}}$ maintains a set of independent modality-specific tokenizers, embedding tables, and prediction heads to handle various modalities. This modular design allows $\text{M}^{\text{3}}$ to process any combination of observation and action modalities.
In addition, $\text{M}^{\text{3}}$ combines several recent improvements from the literature. Specifically, the control policy is intrinsically motivated to reduce the epistemic uncertainty of the world model (Shyam et al., 2019; Sekar et al., 2020), a world model replay buffer further improves sample efficiency (Kauvar et al., 2023), and continuous values are predicted via classification models (Farebrother et al., 2024; Hafner et al., 2023).
@@
To validate the effectiveness of our method, we conducted extensive empirical evaluations across three diverse benchmarks, ranging from the visual Atari 100K, to the continuous proprioception tasks of the DeepMind Control Suite, to Craftax, which combines symbolic 2D grid maps with continuous state features. There, $\text{M}^{\text{3}}$ achieves state-of-the-art sample-efficiency for planning-free world models. Ablation studies further show the contribution of each component.
@@
  • We propose $\text{M}^{\text{3}}$, a widely applicable token-based world model agent that follows a modular framework.
  • We conduct extensive empirical evaluation, providing strong evidence for the effectiveness of TBWMs as general methods. There, $\text{M}^{\text{3}}$ achieves state-of-the-art (planning-free) performance across all benchmarks.
  • To facilitate future research and broader adoption of TBWMs, we open-source our code and trained models.
@@ 2 Method
$\text{M}^{\text{3}}$ builds on REM (Cohen et al., 2024). The agent comprises a representation module $\mathcal{V}$, a world model $\mathcal{M}$, a controller $\mathcal{C}$, and a replay buffer. To facilitate a modular design, following REM, each module is optimized separately. The training process of the agent involves a repeated cycle of four steps: data collection, representation learning ($\mathcal{V}$), world model learning ($\mathcal{M}$), and control learning in imagination ($\mathcal{C}$).
@@ 2.1 The Representation Module $\mathcal{V}$
$\mathcal{V}$ is responsible for encoding and decoding raw observations and actions. It is a modular tokenization system with encoder-decoder pairs for different input modalities. Encoders produce fixed-length token sequences, creating a common interface that enables combining tokens from various sources into a unified representation. After embedding, these token sequences are concatenated into a single representation, as described in Section 2.2. Note that encoder-decoder pairs need not be learning-based methods. Although learned pairs are optimized independently. This design enables $\mathcal{V}$ to deal with any combination of input modalities, provided the respective encoder-decoder pairs.
@@
$\text{M}^{\text{3}}$ natively supports four modalities: images, continuous vectors, categorical variables, and image-like multi-channel grids of categorical variables, referred to as ”2D categoricals”. More formally, 2D categoricals are elements of $([k_{1}]\times[k_{2}]\times\ldots\times[k_{C}])^{m\times n}$ where $k_{1},\ldots,k_{C}$ are per channel vocabulary sizes, $C$ is the number of channels, $m,n$ are spatial dimensions, and $[k]=\{1,\ldots,k\}$.
Following REM, we use a VQ-VAE (Esser et al., 2021; van den Oord et al., 2017) for image observations. For the tokenization of continuous vectors, each feature is quantized to produce a token, as in (Reed et al., 2022). Unbounded vectors are first transformed using the symlog function (Hafner et al., 2023), defined as $\operatorname{symlog}(x)=\operatorname{sign}(x)\ln(1+|x|)$, which compresses the magnitude of large absolute values. Lastly, while no special tokenization is required for categorical inputs, 2D categoricals are flattened along the spatial dimensions to form a sequence of categorical vectors. The embedding of each token vector is obtained by averaging the embeddings of its entries.
@@ 2.2 The World Model $\mathcal{M}$
$\mathcal{M}$ comprises a sequence model $f_{\theta}$ and multiple heads for the prediction of tokens of different observation modalities, rewards, termination signals, and for the estimation of model uncertainty. Concretely, $f_{\theta}$ is a retentive network (RetNet) (Sun et al., 2023) augmented with a parallel observation prediction (POP) (Cohen et al., 2024) mechanism. All heads are implemented as multilayer perceptrons (MLP) with a single hidden layer. We defer the details about these architectures to Appendix A.3.
(a) Tokenization and embedding.
(b) Observation prediction.
@@
Observation-action block: ${\mathbf{X}}_{t}=({\mathbf{X}}^{\text{o}}_{t},{\mathbf{X}}^{\text{a}}_{t})$
Observation block: ${\mathbf{X}}^{\text{o}}_{t}=({\mathbf{X}}^{(1)}_{t},\ldots,{\mathbf{X}}^{(|\kappa|)}_{t})$
@@
$(\mathbf{S}_{t},\mathbf{Y}_{t})=f_{\theta}(\mathbf{S}_{t-1},{\mathbf{X}}_{t})$
@@
$(\cdot,\mathbf{Y}_{t+1}^{\text{u}})=f_{\theta}(\mathbf{S}_{t},{\mathbf{X}}^{\text{u}})$
@@
The world model $\mathcal{M}$ serves as the cornerstone of the entire system. Any controller operating within a world model framework can only perform as well as the underlying world model allows, making its quality a fundamental limiting factor. In deep learning methods, the model’s performance depends heavily on the quality of its training data. Accurate dynamics modeling requires comprehensive data collection that captures the full spectrum of possible environmental behaviors. This presents a particular challenge in online RL, where the controller must systematically and efficiently explore its environment. Success depends on intelligently guiding the controller toward unexplored or undersampled regions of the dynamics space. An effective approach to this challenge involves estimating the world model’s epistemic uncertainty and directing the controller to gather data from regions where this uncertainty is highest (Schmidhuber, 2010; Shyam et al., 2019; Sekar et al., 2020).
Our approach estimates epistemic uncertainty using an ensemble of $N_{\text{ens}}=4$ next observation prediction heads $\left\{p_{\phi_{i}}(\hat{\mathbf{z}}|\operatorname{sg}(\mathbf{Y}^{\text{u}}))\right\}_{i=1}^{N_{\text{ens}}}$ with parameters $\{\phi_{i}\}_{i=1}^{N_{\text{ens}}}$ (Sekar et al., 2020; Lakshminarayanan et al., 2017) where $\operatorname{sg}(\cdot)$ is the stop gradient operator. To quantify disagreement between the ensemble’s distributions, we employ the Jensen-Shannon divergence (JSD) (Shyam et al., 2019). For probability distributions $P_{1},\ldots,P_{n}$, the JSD is defined as:
$\operatorname{JSD}(P_{1},\ldots,P_{n})=\mathcal{H}(\frac{1}{n}\sum_{i=1}^{n}P_{i})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{H}(P_{i})$
@@
$\bar{r}_{t}=w^{\text{int}}r^{\text{int}}_{t}+w^{\text{ext}}\hat{r}_{t}$
@@
Recent work has demonstrated that prioritizing replay buffer sampling during world model training could lead to significant performance gains in intrinsically motivated agents (Kauvar et al., 2023). While their approach showed promise, it required extensive hyperparameter tuning in practice. We propose a simpler, more robust prioritization scheme for world model training.
Here, the replay buffer maintains a world model loss value for each stored example, with newly added examples assigned a high initial loss value of $10$. During $\mathcal{M}$’s training, we sample each batch using a mixture of uniform and prioritized sampling, controlled by a single parameter $\alpha\in[0,1]$ that determines the fraction of prioritized samples. For the prioritized portion, we sample examples proportional to their softmax-transformed losses $p_{i}=\operatorname*{softmax}(\mathcal{L})_{i}$. The loss values are updated after each world model optimization step using the examples’ current batch losses.
@@ 2.3 The Controller $\mathcal{C}$
$\mathcal{C}$ is an extended version of the actor-critic of REM (Cohen et al., 2024) that supports additional observation and action spaces and integrates a classification-based value prediction.
@@
At the core of $\mathcal{C}$’s architecture, parameterized by $\psi$, is an LSTM (Hochreiter & Schmidhuber, 1997) sequence model. At each step $t$, upon observing $\mathbf{z}_{t}$, a set of modality-specific encoders map each modality tokens $\mathbf{z}^{(i)}_{t}$ to a latent vector $\mathbf{x}^{(i)}$, where we abuse our notation $\mathbf{x}$ as the context of the discussion is limited to $\mathcal{C}$. The latents are then fused by a fully-connected network to obtain a single vector $\mathbf{x}=g_{\psi}(\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(|\kappa|)})$. $\mathbf{x}_{t}\in\mathbb{R}^{d_{\mathcal{C}}}$ is processed by $\mathcal{C}$’s sequence model to produce $\mathbf{h}_{t},\mathbf{c}_{t}=\operatorname{LSTM}(\mathbf{x}_{t},\mathbf{h}_{t-1},\mathbf{c}_{t-1};\psi)$ where $\mathbf{h}_{t},\mathbf{c}_{t}$ are the LSTM’s hidden and cell states, respectively. Lastly, two linear output layers produce the logits from which the actor and critic outputs $\pi(\mathbf{a}_{t}|\mathbf{h}_{t}),\hat{V}^{\pi}(\mathbf{h}_{t})$ are derived. For continuous action spaces, the actor uses a categorical distribution over a uniformly spaced discrete subset of $[-1,1]$. We defer the full details about the encoding process to Appendix A.4.
Classification Based Prediction of Continuous Values
Robustly handling unbounded reward signals has long been challenging as they can vary dramatically in both magnitude and frequency. (Hafner et al., 2023) addressed this challenge by using a classification network that predicts the weights of exponentially spaced bins and employed a two-hot loss for the network’s optimization. (Farebrother et al., 2024) studied the use of cross-entropy loss in place of the traditional mean squared error loss for value-based deep RL methods. In their work, the HL-Gauss method was shown to significantly outperform the two-hot loss method. Following these developments, we adopt (Hafner et al., 2023)’s approach of using a classification network with exponential bins, while implementing the HL-Gauss method for the optimization of the network. Concretely, the critic’s value estimates are predicted using a linear output layer parameterized by $\mathbf{W}\in\mathbb{R}^{m\times d_{\mathcal{C}}}$ with $m=128$ outputs corresponding to $m$ uniform bins defined by $m+1$ endpoints $\mathbf{b}=(b_{0},\ldots,b_{m})$. The predicted value is given by
@@
$\mathcal{C}$ is trained entirely from simulated experience generated through interaction with $\mathcal{M}$. Specifically, $\mathcal{M}$ and $\mathcal{C}$ are initialized with a short trajectory segment sampled uniformly from the replay buffer and interact for $H$ steps. An illustration of this process is given in Figure 3 (orange path). $\lambda$-returns are computed for each generated trajectory segment and are used as targets for critic learning. For policy learning, a REINFORCE (Sutton et al., 1999) objective is used, with a $\hat{V}^{\pi}$ baseline for variance reduction. See Appendix A.4.2 for further details.
Game Random Human DreamerV3 TWM STORM DIAMOND REM $\text{M}^{\text{3}}$ (ours)
Alien 227.8 7127.7 959.4 674.6 983.6 744.1 607.2 687.2
Amidar 5.8 1719.5 139.1 121.8 204.8 225.8 95.3 102.4
Assault 222.4 742.0 705.6 682.6 801.0 1526.4 1764.2 1822.8
Asterix 210.0 8503.3 932.5 1116.6 1028.0 3698.5 1637.5 1369.1
BankHeist 14.2 753.1 648.7 466.7 641.2 19.7 19.2 347.1
Boxing 0.1 12.1 78.0 77.5 79.7 86.9 87.5 93.5
Breakout 1.7 30.5 31.1 20.0 15.9 132.5 90.7 148.9
DemonAttack 152.1 1971.0 303.3 350.2 164.6 288.1 5738.6 4787.6
Freeway 0.0 29.6 0.0 24.3 0.0 33.3 32.3 31.9
Frostbite 65.2 4334.7 909.4 1475.6 1316.0 274.1 240.5 258.4
Gopher 257.6 2412.5 3730.0 1674.8 8239.6 5897.9 5452.4 4363.2
Hero 1027.0 30826.4 11160.5 7254.0 11044.3 5621.8 6484.8 7466.8
Jamesbond 29.0 302.8 444.6 362.4 509.0 427.4 391.2 678.0
Kangaroo 52.0 3035.0 4098.3 1240.0 4208.0 5382.2 467.6 6656.0
Krull 1598.0 2665.5 7781.5 6349.2 8412.6 8610.1 4017.7 6677.3
MsPacman 307.3 6951.6 1326.9 1588.4 2673.5 1958.2 962.5 1282.7
Pong -20.7 14.6 18.4 18.8 11.3 20.4 18.0 19.9
PrivateEye 24.9 69571.3 881.6 86.6 7781.0 114.3 99.6 100.0
Qbert 163.9 13455.0 3405.1 3330.8 4522.5 4499.3 743.0 2425.6
Seaquest 68.4 42054.7 618.0 774.4 525.2 551.2 1036.7 1800.4
UpNDown 533.4 11693.2 7567.1 15981.7 7985.0 3856.3 3757.6 10416.5
#Superhuman (↑) 0 N/A 9 8 9 11 12 13
Mean (↑) 0.000 1.000 1.124 0.956 1.222 1.459 1.222 1.645
Median (↑) 0.000 1.000 0.485 0.505 0.425 0.373 0.280 0.982
IQM (↑) 0.000 1.000 0.487 0.459 0.561 0.641 0.673 0.990
Optimality Gap (↓) 1.000 0.000 0.510 0.513 0.472 0.480 0.482 0.412

### v3

# Uncovering Untapped Potential in Sample-Efficient World Model Agents

@@ Abstract
World model (WM) agents enable sample-efficient reinforcement learning by learning policies entirely from simulated experience. However, existing token-based world models (TBWMs) are limited to visual inputs and discrete actions, restricting their adoption and applicability. Moreover, although both intrinsic motivation and prioritized WM replay have shown promise in improving WM performance and generalization, they remain underexplored in this setting, particularly in combination. We introduce Simulus, a highly modular TBWM agent that integrates (1) a modular multi-modality tokenization framework, (2) intrinsic motivation, (3) prioritized WM replay, and (4) regression-as-classification for reward and return prediction. Simulus achieves state-of-the-art sample efficiency for planning-free WMs across three diverse benchmarks. Ablation studies reveal the individual contribution of each component while highlighting their synergy. Our code and model weights are publicly available at https://github.com/leor-c/Simulus.
@@ 1 Introduction
Sample efficiency refers to the ability of a reinforcement learning (RL) algorithm to learn effective policies using as few environment interactions as possible. In many real-world domains such as robotics, autonomous driving, and healthcare, this is particularly critical, as interactions are costly, slow, or constrained. World model agents, methods that learn control entirely from simulated experience generated by a learned dynamics model, have emerged as a promising approach to improving sample efficiency [17, 20, 37, 57].
@@
Here, we focus on token-based world models (TBWMs) [37, 38, 11], sample-efficient RL methods that learn the dynamics entirely within a learned discrete token space, where each observation comprises a sequence of tokens. Evidently, most large scale world models [1, 12, 14] operate on multi-token observations, suggesting that such representations are advantageous at scale. TBWMs offer a clear modular design, separating the optimization of its representation, dynamics, and control models. As modular systems, TBWMs are easier to scale, develop, study, and deploy, as individual modules can be treated independently without interfering and are easier to master through divide and conquer. In addition, such separation leads to simpler optimization objectives and avoids interference between them (Appendix B.1).
However, existing TBWMs are restricted to image observations and discrete actions, such as Atari games, limiting both their adoption and broader applicability, as their effectiveness in diverse environments and modalities remains unclear. While multi-modality tokenization approaches exist for large-scale offline settings [43, 47, 33, 34], these methods rely on large vocabularies (e.g., 33K tokens) which are inefficient for online, data-limited regimes. Whether substantially smaller vocabularies can preserve competitive precision and performance is still an open question, leaving these approaches underexplored for sample-efficient RL.
Furthermore, despite compelling results [46, 49, 30], intrinsic motivation and prioritized world model replay remain underexplored in sample-efficient world model agents [19, 20, 60, 11, 3], particularly in combination. We conjecture that intrinsic motivation is underused as it may steer the agent toward task-irrelevant regions, potentially wasting limited interaction budget. Prioritized replay [30], while promising, lacks robust empirical support and proved challenging to tune in our experiments.
To address these limitations, we propose Simulus, a modular world model that extends a recent TBWM method [11] by integrating several powerful advances from the literature: (1) a modular multi-modality tokenization framework for handling arbitrary combinations of observation and action modalities, (2) intrinsic motivation for epistemic uncertainty reduction [50, 49], (3) prioritized world model replay [30], and (4) regression-as-classification (RaC) for reward and return prediction [16, 20].
@@
To evaluate the impact of the proposed components, we conducted extensive empirical evaluations across three diverse benchmarks, ranging from the visual Atari 100K [28], to the continuous proprioception tasks of the DeepMind Control Suite [53], to Craftax [35], which combines symbolic 2D grid maps with continuous state features. There, Simulus achieves state-of-the-art sample-efficiency for planning-free world models. Ablation studies further show the contribution and effectiveness of each component.
@@
  • Through extensive empirical evaluations across three diverse benchmarks, we show that intrinsic motivation, prioritized replay, and RaC significantly improve performance in the sample-efficiency setting for world model agents, particularly when combined.
  • Our results demonstrate the effectiveness of the multi-modality tokenization approach, as Simulus achieves state-of-the-art planning-free performance across benchmarks, establishing TBWMs as widely-applicable methods.
  • We propose Simulus, a versatile TBWM agent that follows a highly modular design, offering a solid foundation for future developments. To support future research and adoption of TBWMs, we open-source our code and weights.
Figure 1: Results overview. Simulus exhibits state-of-the-art sample-efficiency performance for planning-free methods across all three benchmarks. $\dagger$ [44], $\ddagger$ [13].
@@ 2 Method
Simulus builds on REM [11]. The agent comprises a representation module $\mathcal{V}$, a world model $\mathcal{M}$, a controller $\mathcal{C}$, and a replay buffer. To facilitate a modular design, following REM, each module is optimized separately. The training process of the agent involves a repeated cycle of four steps: data collection, representation learning ($\mathcal{V}$), world model learning ($\mathcal{M}$), and control learning in imagination ($\mathcal{C}$).
@@ 2.1 The Representation Module $\mathcal{V}$
$\mathcal{V}$ is responsible for encoding and decoding raw observations and actions. It is a modular tokenization system with encoder-decoder pairs for different input modalities. Encoders produce fixed-length token sequences, creating a common interface that enables combining tokens from various sources into a unified representation. After embedding, these token sequences are concatenated into a single representation, as described in Section 2.2. Note that encoder-decoder pairs need not be learning-based methods; however, when learned, they are optimized independently. This design enables $\mathcal{V}$ to deal with any combination of input modalities, provided the respective encoder-decoder pairs.
@@
Simulus natively supports four modalities: images, continuous vectors, categorical variables, and image-like multi-channel grids of categorical variables, referred to as "2D categoricals". More formally, 2D categoricals are elements of $([k_{1}]\times[k_{2}]\times\ldots\times[k_{C}])^{m\times n}$ where $k_{1},\ldots,k_{C}$ are per channel vocabulary sizes, $C$ is the number of channels, $m,n$ are spatial dimensions, and $[k]=\{1,\ldots,k\}$.
Following REM, we use a VQ-VAE [15, 55] for image observations. For the tokenization of continuous vectors, each feature is quantized to produce a token, as in [43]. However, to improve learning efficiency, we reduce the vocabulary size from 1024 to 125, and modify the quantization levels for optimal coverage (Appendix A.2.2). Unbounded vectors are first transformed using the symlog function [20], defined as $\operatorname{symlog}(x)=\operatorname{sign}(x)\ln(1+|x|)$, which compresses the magnitude of large absolute values. Lastly, while no special tokenization is required for categorical inputs, 2D categoricals are flattened along the spatial dimensions to form a sequence of categorical vectors. The embedding of each token vector is obtained by averaging the embeddings of its entries.
@@ 2.2 The World Model $\mathcal{M}$
$\mathcal{M}$ comprises a sequence model $f_{\theta}$ and multiple heads for the prediction of tokens of different observation modalities, rewards, termination signals, and for the estimation of model uncertainty. Concretely, $f_{\theta}$ is a retentive network (RetNet) [51] augmented with a parallel observation prediction (POP) [11] mechanism. All heads are implemented as multilayer perceptrons (MLP) with a single hidden layer. We defer the details about these architectures to Appendix A.3.
@@
Observation-action block: ${\mathbf{X}}_{t}=({\mathbf{X}}^{\text{o}}_{t},{\mathbf{X}}^{\text{a}}_{t}),$
Observation block: ${\mathbf{X}}^{\text{o}}_{t}=({\mathbf{X}}^{(1)}_{t},\ldots,{\mathbf{X}}^{(|\kappa|)}_{t}),$
@@
(a) Observation tokenization and embedding.
(b) Prediction of observation tokens.
@@
$(\mathbf{S}_{t},\mathbf{Y}_{t})=f_{\theta}(\mathbf{S}_{t-1},{\mathbf{X}}_{t}),$
@@
$(\cdot,\mathbf{Y}_{t+1}^{\text{u}})=f_{\theta}(\mathbf{S}_{t},{\mathbf{X}}^{\text{u}}),$
@@
The world model $\mathcal{M}$ serves as the cornerstone of the entire system. Any controller operating within a world model framework can only perform as well as the underlying world model allows, making its quality a fundamental limiting factor. In deep learning methods, the model’s performance depends heavily on the quality of its training data. Accurate dynamics modeling requires comprehensive data collection that captures the full spectrum of possible environmental behaviors. This presents a particular challenge in online RL, where the controller must systematically and efficiently explore its environment. Success depends on intelligently guiding the controller toward unexplored or undersampled regions of the dynamics space. An effective approach to this challenge involves estimating the world model’s epistemic uncertainty and directing the controller to gather data from regions where this uncertainty is highest [46, 50, 49].
Our approach estimates epistemic uncertainty using an ensemble of $N_{\text{ens}}=4$ next observation prediction heads $\left\{p_{\phi_{i}}(\hat{\mathbf{z}}|\operatorname{sg}(\mathbf{Y}^{\text{u}}))\right\}_{i=1}^{N_{\text{ens}}}$ with parameters $\{\phi_{i}\}_{i=1}^{N_{\text{ens}}}$ [49, 31] where $\operatorname{sg}(\cdot)$ is the stop gradient operator. To quantify disagreement between the ensemble’s distributions, we employ the Jensen-Shannon divergence (JSD) [50]. For probability distributions $P_{1},\ldots,P_{n}$, the JSD is defined as:
$\operatorname{JSD}(P_{1},\ldots,P_{n})=\mathcal{H}(\frac{1}{n}\sum_{i=1}^{n}P_{i})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{H}(P_{i}),$
@@
$\bar{r}_{t}=w^{\text{int}}r^{\text{int}}_{t}+w^{\text{ext}}\hat{r}_{t},$
@@
Recent work has demonstrated that prioritizing replay buffer sampling during world model training could lead to significant performance gains in intrinsically motivated agents [30]. While their approach showed promise, it required extensive hyperparameter tuning in practice. We propose a simpler, more robust prioritization scheme for world model training.
Here, the replay buffer maintains a world model loss value for each stored example, with newly added examples assigned a high initial loss value $\nu_{0}$. During $\mathcal{M}$’s training, we sample each batch using a mixture of uniform and prioritized sampling, controlled by a single parameter $\alpha\in[0,1]$ that determines the fraction of prioritized samples. For the prioritized portion, we sample examples proportional to their softmax-transformed losses $p_{i}=\operatorname*{softmax}(\mathcal{L})_{i}$. The loss values are updated after each world model optimization step using the examples’ current batch losses.
@@ 2.3 The Controller $\mathcal{C}$
$\mathcal{C}$ is an extended version of the actor-critic of REM [11] that supports additional observation and action spaces and implements regression-as-classification for return predictions.
@@
At the core of $\mathcal{C}$’s architecture, parameterized by $\psi$, is an LSTM [26] sequence model. At each step $t$, upon observing $\mathbf{z}_{t}$, a set of modality-specific encoders map each modality tokens $\mathbf{z}^{(i)}_{t}$ to a latent vector $\mathbf{x}^{(i)}$, where we abuse our notation $\mathbf{x}$ as the context of the discussion is limited to $\mathcal{C}$. The latents are then fused by a fully-connected network to obtain a single vector $\mathbf{x}=g_{\psi}(\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(|\kappa|)})$. $\mathbf{x}_{t}\in\mathbb{R}^{d_{\mathcal{C}}}$ is processed by $\mathcal{C}$’s sequence model to produce $\mathbf{h}_{t},\mathbf{c}_{t}=\operatorname{LSTM}(\mathbf{x}_{t},\mathbf{h}_{t-1},\mathbf{c}_{t-1};\psi)$ where $\mathbf{h}_{t},\mathbf{c}_{t}$ are the LSTM’s hidden and cell states, respectively. Lastly, two linear output layers produce the logits from which the actor and critic outputs $\pi(\mathbf{a}_{t}|\mathbf{h}_{t}),\hat{V}^{\pi}(\mathbf{h}_{t})$ are derived. For continuous action spaces, the actor uses a categorical distribution over a uniformly spaced discrete subset of $[-1,1]$. We defer the full details about the encoding process to Appendix A.4.
Regression as Classification for Reward and Return Prediction
Robustly handling unbounded reward signals has long been challenging as they can vary dramatically in both magnitude and frequency. Hafner et al. [20] addressed this challenge by using a classification network that predicts the weights of exponentially spaced bins and employed a two-hot loss for the network’s optimization. Farebrother et al. [16] studied the use of cross-entropy loss in place of the traditional mean squared error loss for value-based deep RL methods. In their work, the HL-Gauss method was shown to significantly outperform the two-hot loss method. Building on these developments, we adopt the classification network with exponential bins from [20], and apply the HL-Gauss method for its optimization. Concretely, the critic’s value estimates are predicted using a linear output layer parameterized by $\mathbf{W}\in\mathbb{R}^{m\times d_{\mathcal{C}}}$ with $m=128$ outputs corresponding to $m$ uniform bins defined by $m+1$ endpoints $\mathbf{b}=(b_{0},\ldots,b_{m})$. The predicted value is given by
@@
$\mathcal{C}$ is trained entirely from simulated experience generated through interaction with $\mathcal{M}$. Specifically, $\mathcal{M}$ and $\mathcal{C}$ are initialized with a short trajectory segment sampled uniformly from the replay buffer and interact for $H$ steps. An illustration of this process is given in Figure 3 (orange path). $\lambda$-returns are computed for each generated trajectory segment and are used as targets for critic learning. For policy learning, a REINFORCE [52] objective is used, with a $\hat{V}^{\pi}$ baseline for variance reduction. See Appendix A.4.2 for further details.